# thanks: Corresponding author

### v1

@@ V Conclution
The 21-cm forest signal primarily comes from the neutral hydrogen within and around halos, so we focus on small-scale modeling. However, modeling at larger scales is also important. For high-redshift radio point sources, the neutral segments in their light paths are numerous and long, making 2 Mpc scale modeling inadequate. For low-redshift radio point sources, modeling at larger scales requires considering the ionized fraction of neutral hydrogen, which helps explore the evolution of the reionization history.
@@ Acknowledgments
We thank Yichao Li and Yidong Xu for helpful discussions. This work was supported by the National SKA Program of China (grant Nos. 2022SKA0110200 and 2022SKA0110203), the National Natural Science Foundation of China (grant Nos. 12473001, 11975072, 11875102, and 11835009), and the National 111 Project (Grant No. B16009).

### v2

@@ V Conclusion
The 21-cm forest signal primarily comes from the neutral hydrogen within and around halos, so we focus on small-scale modeling. However, modeling at larger scales is also important. For high-redshift radio point sources, the neutral segments in their light paths are numerous and long, making $2{\rm~Mpc}$ scale modeling inadequate. For low-redshift radio point sources, modeling at larger scales requires considering the ionized fraction of neutral hydrogen, which helps explore the evolution of the reionization history.
@@ Acknowledgments
We would like to thank Yidong Xu for the early-stage suggestion to use the halo model for analytical modeling, for the insightful discussions regarding the halo model method, and for providing the code about small-scale simulation of 21-cm forest. Additionally, we also thank Yichao Li and Qi Niu for the fruitful discussions. This work was supported by the National SKA Program of China (Grant No. 2022SKA0110200 and No. 2022SKA0110203), the National Natural Science Foundation of China (Grant No. 12533001, No. 12473001, No. 11975072, No.11875102, and No. 11835009), and the National 111 Project (Grant No. B16009).
Data Availability
The data are not publicly available. The data are available from the authors upon reasonable request.